A. Find the area underneath the given curve.


1.) y= -2x² + 8

from x=0 to x=1


2.) y= -x² + 7

from x=0 to x=1


B. Find the area enclosed by the given curve, the x-axis, and the given lines.


1.) y= -1/3x² + 10

from x=1 and x=3


2.) y= x³ - 8

from x= -1 to x =2


C. Find the area bounded by the given curve and line.


1.) y= x² + 3 and y= 7


2.) y= 2x² and y= 4x + 6


3.) y= -x² + 4x and y= x²


4.) y= x³ -6x² + 8x and y= x² -4x


D. Find the area bounded by y=2; y=√x+2; and y=2 - x.


1
Expert's answer
2021-03-31T14:05:50-0400

A)

1)

"\intop_0^1(2X^2+8)\\
=8.66"

2)

"\int_0^1-x^2+7\\={20\over3}=6.667"


B)

1)

"y={-1\over 3}x^2+10 \\x=1\\x=3 \\
\therefore\int_1^3({-1\over 3}x^2 +10)
\\\boxed{=17.11}"



2)

"y_{1}=x^{3}-8\\
x=1\\
x=2
\\
\therefore 
\int_1^2(x^3-8x)
\\\boxed{=-4.5}"



C]

1)

"y_{1}=x^{2}+3
\\y=7
\\\therefore\int_{-2}^{2}\left(7-x^{2}+3\right)dx\\\boxed{=34.667}"


2)

"y_1=2x^2\\
y_2=4x+6\\
\int_{-1}^{3}\left(y_{1}-y_{2}\right)dx
\\=
-21.33"


3)

"y_{1}=x^{2}\\y_{2}=-x^{2}+4x\\\int_{0}^{2}\left(y_{2}-y_{1}\right)dx\\=2.66"


4)

"y_{1}=x^{3}-6x^{2}+8x\\y_{2}=x^{2}-4x\\\int_{0}^{3}\left(y_{1}-y_{2}\right)dx+\int_{3}^{4}\left(y_{2}-y_{1}\right)dx\\\boxed{=11.833}"


D}


"y=2\\
y_{2}=\sqrt{x}+2\\y_{3}=2-x"



these curves has no area enclosed so area=0